OpenStudy (anonymous):

factor completely x^3 +3x^2 -36x -108

OpenStudy (xkehaulanix):

\[x ^{3} + 3x ^{2} - 36x - 108\]\[=x ^{3} - 36x + 3x ^{2} - 108\]\[= x (x ^{2}-36)+3(x ^{2}-36)\]\[=(x ^{2}-36)(x+3)\]\[=(x+6)(x-6)(x+3)\]